OpenStudy (anonymous):

If f(x) = 0.5(1 – x), what is the value of f(–1)?

OpenStudy (ajprincess):

plug in -1 in place of x in f(x)=0.5(1 – x). Does that help?

OpenStudy (anonymous):

yes but no

OpenStudy (ajprincess):

f(x)=0.5(1 – x) f(-1)=0.5(1-(-1))

OpenStudy (anonymous):

so what is do is, 0.5 (1-(-1)), i do that math problem to find the answer?

OpenStudy (ajprincess):

0.5 (1-(-1))=?

OpenStudy (anonymous):

i dont understand how to work it out

OpenStudy (ajprincess):

what is 1-(-1)?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

sorry, it is 2 because two negatives makes a positive so it would be 2, am i correct

OpenStudy (ajprincess):

yup u r right:) nw multiply 2 by 0.5. 2*0.5=?

OpenStudy (anonymous):

1 so 1 would be my answer

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

ok thank u can i ask u more questions

OpenStudy (ajprincess):

ya sure

OpenStudy (anonymous):

ok, thank you

OpenStudy (anonymous):

how do you find the volume of a cylinder

OpenStudy (ajprincess):

use the formula  Volume of the cylinder=\(\large\pi r^2l\)